Brief Exercise 22-09 Your answer is partially correct. Try again Ivanhoe Industries expects credit sales for January, FebruarPlease correct problem.
Answer #1
January February March
January 156,000 52,000
February (261,900*75)% = 196,425 (261,900*25%)= 65,475
March (316,500*75%)= 237,375
156,000 248,428 302,850

Comment if you face any issues